# Vireo: Leveraging Depth and Language for Open-Vocabulary Domain-Generalized Semantic Segmentation

### v2

@@ Abstract
Open-Vocabulary semantic segmentation (OVSS) and domain generalization in semantic segmentation (DGSS) highlight a subtle complementarity that motivates Open-Vocabulary Domain-Generalized Semantic Segmentation (OV-DGSS). OV-DGSS aims to generate pixel-level masks for unseen categories while maintaining robustness across unseen domains, a critical capability for real-world scenarios such as autonomous driving in adverse conditions. We introduce Vireo, a novel single-stage framework for OV-DGSS that unifies the strengths of OVSS and DGSS for the first time. Vireo builds upon the frozen Visual Foundation Models (VFMs) and incorporates scene geometry via Depth VFMs to extract domain-invariant structural features. To bridge the gap between visual and textual modalities under domain shift, we propose three key components: (1) GeoText Prompts, which align geometric features with language cues and progressively refine VFM encoder representations; (2) Coarse Mask Prior Embedding (CMPE) for enhancing gradient flow for faster convergence and stronger textual influence; and (3) the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH), which fuses refined structural and semantic features for robust prediction. Comprehensive evaluation on these components demonstrates the effectiveness of our designs. Our proposed Vireo achieves the state-of-the-art performance and surpasses existing methods by a large margin in both domain generalization and open-vocabulary recognition, offering a unified and scalable solution for robust visual understanding in diverse and dynamic environments. Code is available at https://github.com/anonymouse-9c53tp182bvz/Vireo.
@@ 1 Introduction
Figure 1: Overview of Vireo and effectiveness. (I) Performance comparison under various adverse conditions across ACDC and DELIVER datasets, showing that Vireo consistently outperforms existing methods. (II) Qualitative visualization of segmentation predictions (left) and attention maps (right) under extreme night scenes, illustrating Vireo’s robustness and precise alignment with semantic cues. (III) Architectural comparison: (a) Traditional OVSS and (b) DGSS pipelines freeze or fine-tune VFM separately without cross-modal integration; (c) Our proposed Vireo introduces GeoText Prompts and Depth-VFM integration to enhance both semantic alignment and domain robustness.
@@
Moreover, in cross-domain scenarios, depth and geometric cues are largely insensitive to variations in illumination and texture[19, 20]. They supply reliable spatial constraints, easing the distribution shift of RGB features and sharpening boundary localization. Recent studies such as DepthForge[21] have shown that injecting depth prompts into a frozen VFM boosts domain generalization. Motivated by these findings, we adopt DepthAnything V2 as our Depth VFM: its diverse pre-training delivers consistent depth estimation across domains, and keeping it fully frozen incurs minimal training cost while preserving real-time inference speed.
In this paper, we propose a VFM-based single-stage pipeline for OV-DGSS, termed Vireo. Specifically, in the encoder, both the VFM and DepthAnything modules are kept frozen. The VFM is leveraged to robustly capture cross-domain visual features, while DepthAnything extracts the scene’s intrinsic geometric structure. On this basis, we introduce Geometric-Text Prompts (GeoText Prompts) to fuse the extracted structural features with manually provided textual cues, progressively refining the feature maps between the frozen VFM layers. To mitigate the slow convergence caused by sparse gradients propagating through the frozen encoder, we introduce a Coarse Mask Prior Embedding (CMPE) at the beginning of the decoder to inject denser gradient signals. This design accelerates the convergence of mask supervision and further strengthens the influence of textual priors. Subsequently, we design the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH) to strengthen the synergistic integration of structural and textual modalities, ensuring that both cross-domain structural features and open-vocabulary cues learned by the GeoText prompts are fully utilized in the final prediction.
For challenge 1: We find that GeoText Prompts not only capture structural and semantic cues within the frozen VFM encoder but also guide the progressive refinement of its feature representations. For challenge 2: The CMPE enhances gradient back-propagation into the encoder, while the redesigned DOV-VEH deepens the fusion of visual and textual priors. These components form a unified framework that simultaneously achieves domain robustness and strong open-vocabulary generalization. Our main contributions are summarized as follows:
@@
(2) We introduce GeoText Prompts to progressively refine frozen VFM features by injecting geometric cues from DepthAnything and aligning them with textual semantics, enabling structural-semantic fusion across encoder layers.
@@
Figure 2: Overview of the proposed Vireo framework for OV-DGSS. (a) Framework Overview. Geometric-Text Prompts permeate our model: they’re injected in Tunable Vireo to align domain priors, reused in CMPE to guide multi-scale feature fusion, and employed in DOV-VEH as queries for open-vocabulary segmentation, forming a unified end-to-end loop. (b) Tunable Vireo. GeoText Prompts are used to inject structural-textual priors across multiple layers. Depth-aware fusion and attention optimization are applied between intermediate frozen layers to progressively refine representations. (c) Coarse Mask Prior Embedding (CMPE). Multi-scale features from VFM are combined with GeoText Prompts to provide dense supervision and gradient signals for downstream modules. (d) Domain-Open-Vocabulary Head (DOV-VEH). Multi-level features are processed by a Pixel Decoder and Transformer Decoder to produce final predictions guided by open-vocabulary text queries.
@@ 3.1 Overview
As illustrated in Figure 2, our framework consists of three key modules tailored for the OV-DGSS:
  • Tunable Vireo with GeoText Prompts: Introduces GeoText Prompts to inject and refine both geometric and textual information across layers of a frozen VFM.
  • Coarse Mask Prior Embedding (CMPE): Generates coarse prior masks to guide segmentation and reinforce gradient flow from the decoder back to the frozen encoder.
  • Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH): Integrates visual, geometric, and semantic features to produce final OV-DGSS predictions.
@@
The input RGB image $I\in\mathbb{R}^{H\times W\times 3}$ is duplicated and independently fed into two frozen encoders: a VFM encoder $\mathcal{F}^{V}$ and a DepthAnything encoder $\mathcal{F}^{D}$. These backbones extract a series of multi-scale features $f^{V}_{l}$ and $f^{D}_{l}$ respectively, where $l$ denotes the encoder layer index in $[1,L]$. In parallel, a set of class labels $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ is transformed into natural language prompts $T=\{\texttt{prompt}_{1},\dots,\texttt{prompt}_{K}\}$, which are encoded using a frozen CLIP text encoder $\mathcal{F}^{T}$ to obtain textual embeddings $t_{k}\in\mathbb{R}^{d}$. These features serve as shared semantic priors across the entire framework and are precomputed once at initialization. In the Tunable Vireo module, each layer receives a tuple $(f^{V}_{l},f^{D}_{l},t_{k})$, and applies the GeoText Prompts $P_{l}$ to align and refine the visual features. This is achieved by computing cross-modal attention maps $\mathcal{A}_{l}=\text{Attn}(P_{l},f^{V}_{l},f^{D}_{l},t_{k})$ followed by feature fusion and projection. The refined output $\hat{f^{V}}_{l}$ is then forwarded back to update the VFM’s layer-wise activations.
@@
We select the refined visual features $\{{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder at layers $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$ for downstream decoding. In the Coarse Mask Prior Embedding (CMPE) module, these features are first upsampled to a unified spatial resolution and passed through a channel-spatial attention gating function $\mathcal{G}(\cdot)$. The gated outputs are then fused to form a global coarse feature $f^{M}\in\mathbb{R}^{d}$. The $f^{M}$ is projected and matched with the text embeddings $t_{k}\in\mathbb{R}^{d}$ to produce a coarse class probability map $\mathcal{M}\in\mathbb{R}^{H\times W\times K}$, where $\mathcal{M}$ serves both as a weak supervision signal and as a prior to construct the query embeddings $\mathcal{Q}$. The query prior is added to the GeoText Prompts and forwarded to the final segmentation head. In the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH), the multi-scale features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ are passed through a pixel decoder $\mathcal{D}_{p}(\cdot)$ to enhance spatial representation, followed by a Transformer decoder $\mathcal{D}_{T}(\cdot)$ that leverages positional embedding. The GeoText Prompts serve as learnable queries and interact with both the decoded features and the text embeddings, producing pixel-level mask embeddings $\mathcal{E}_{\text{mask}}(x,y)\in\mathbb{R}^{d}$ and classification embeddings $\mathcal{E}_{\text{cls}}(k)\in\mathbb{R}^{d}$. The final prediction is ^$\hat{\mathcal{M}}(x,y,k)\in\mathbb{R}^{H\times W\times K}$ provides the pixel-wise semantic prediction with both fine-grained detail and open-vocabulary generalization capability.
3.2 Tunable Vireo with GeoText Prompts
@@
During inference, the input image $I\in\mathbb{R}^{H\times W\times 3}$ is simultaneously processed by a frozen visual encoder $\mathcal{F}^{V}$ and a frozen depth encoder $\mathcal{F}^{D}$ (e.g., DepthAnything). For each selected layer $l\in\{1,\dots,L\}$, we obtain the visual feature map $f^{V}_{l}=\mathcal{F}^{V}_{l}(I)$ and the depth feature map $f^{D}_{l}=\mathcal{F}^{D}_{l}(I)$. Each layer of the Tunable Vireo module receives the tuple $(f^{V}_{l},f^{D}_{l},\{t_{k}\})$ along with a layer-specific GeoText Prompt $P_{l}\in\mathbb{R}^{N\times d}$, where $N$ is the number of learnable prompts.
@@ 3.3 Coarse Mask Prior Embedding (CMPE)
Each $f^{\text{class}}_{k}$ is projected into the embedding space as $e^{\text{class}}_{k}\in\mathbb{R}^{d}$, and combined with a set of learnable query vectors $\{q_{j}\}_{j=1}^{N_{q}}$ to produce the query priors:
@@
The final query priors $\{q^{\text{prior}}_{j}\}_{j=1}^{N_{q}}$ are added to the corresponding GeoText Prompts and forwarded into the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH).
@@ 3.4 Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH)
The DOV-VEH module receives the refined multi-scale features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder, where $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$, along with the updated GeoText Prompts $\{P_{l}\}$. These features are first processed by a pixel decoder $\mathcal{D}_{p}(\cdot)$, which leverages multi-scale cross-attention to extract rich spatial context: $f^{\text{pix}}=\mathcal{D}_{p}\left(\hat{f}^{V}_{l_{i}}\right)$. The fused feature $f^{\text{pix}}\in\mathbb{R}^{H\times W\times d}$ is then compressed via a $1\times 1$ convolution and enriched with sinusoidal positional encoding to preserve spatial structure.
The enhanced features are fed into a Transformer Decoder $\mathcal{D}_{T}(\cdot)$, where the GeoText Prompts act as learnable queries. Through stacked layers of self-attention and cross-attention with $f^{\text{pix}}$, the model captures fine-grained visual-semantic alignment, yielding a set of high-resolution mask features $\mathcal{E}_{\text{mask}}(x,y)\in\mathbb{R}^{d}$ at each spatial position.
Simultaneously, the GeoText Prompts $\{P\}$ are passed through a two-layer MLP and then interact with the text embeddings $\{t_{k}\}$ (precomputed from the CLIP text encoder) to produce classification-level representations $\mathcal{E}_{\text{cls}}(k)\in\mathbb{R}^{d}$. The final segmentation prediction ^$\hat{\mathcal{M}}\in\mathbb{R}^{H\times W\times K}$ is generated via an Einstein summation over the two embeddings:
^$\hat{\mathcal{M}}(x,y,k)=\sum_{d=1}^{D}\mathcal{E}_{\text{mask}}(x,y,d)\cdot\mathcal{E}_{\text{cls}}(k,d),$ (5)
@@ 4.4 Ablation Study
Figure 5: Visualizations of attention maps and affinity maps under differnet scenes, where CMPE denotes the Coarse Mask Prior Embedding, GTP represents the GeoText Prompts, respectively.
@@
$+$ concat $f^{d}_{i}$ 70.8 56.1 79.4 72.8 63.6 74.2
$+$ Prompt Depth Anything 70.4 55.5 79.7 72.0 63.8 73.9
$+$ Depth Anything V2 71.5 56.7 80.5 73.3 64.4 74.5
$+$ DA + AO 72.2 57.4 80.9 74.2 65.1 75.0
$+$ DOV-VEH 70.9 56.2 79.8 72.8 63.7 74.2
$+$ CMPE 71.6 56.9 80.2 73.4 64.1 74.6
$+$ GeoText Prompts 74.0 58.4 81.1 74.8 65.3 75.3
@@
Fine-tune Trainable Params mIoU Avg.
@@
Fine-tune Trainable Params mIoU Avg.
@@
Fine-tune Trainable Params mIoU Avg.
@@
Fine-tune Trainable Params mIoU Avg.
@@
In Table 5, using Depth Anything V2 alone yields an approximate $0.6\%$ mIoU gain across all six scenarios, and adding depth augmentation with attention optimization (DA + AO) delivers a further $\approx 1\%$ improvement. The DOV-VEH and CMPE modules each add $0.5\%$$0.8\%$, validating mask vectors and dense gradient embedding. GeoText Prompts on its own provides a substantial $\approx 4.4\%$ boost, underscoring the complementary benefits of fusing semantic and geometric cues.
Pronounced Attention Focus. Figure 5 further confirm that GeoText Prompts steer the model toward geometry-sensitive regions, while CMPE strengthens gradients on foreground masks; compared with baselines, Vireo’s focus is tighter on scene structure and semantic boundaries, explaining its consistent advantage in cross-domain and open-vocabulary settings.

### v3

# Leveraging Depth and Language for Open-Vocabulary Domain-Generalized Semantic Segmentation

@@ Abstract
Open-Vocabulary semantic segmentation (OVSS) and domain generalization in semantic segmentation (DGSS) highlight a subtle complementarity that motivates Open-Vocabulary Domain-Generalized Semantic Segmentation (OV-DGSS). OV-DGSS aims to generate pixel-level masks for unseen categories while maintaining robustness across unseen domains, a critical capability for real-world scenarios such as autonomous driving in adverse conditions. We introduce Vireo, a novel single-stage framework for OV-DGSS that unifies the strengths of OVSS and DGSS for the first time. Vireo builds upon the frozen Visual Foundation Models (VFMs) and incorporates scene geometry via Depth VFMs to extract domain-invariant structural features. To bridge the gap between visual and textual modalities under domain shift, we propose three key components: (1) GeoText Query, which align geometric features with language cues and progressively refine VFM encoder representations; (2) Coarse Mask Prior Embedding (CMPE) for enhancing gradient flow for faster convergence and stronger textual influence; and (3) the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH), which fuses refined structural and semantic features for robust prediction. Comprehensive evaluation on these components demonstrates the effectiveness of our designs. Our proposed Vireo achieves the state-of-the-art performance and surpasses existing methods by a large margin in both domain generalization and open-vocabulary recognition, offering a unified and scalable solution for robust visual understanding in diverse and dynamic environments. Code is available at https://github.com/SY-Ch/Vireo.
@@ 1 Introduction
Figure 1: Overview of Vireo and effectiveness. (I) Performance comparison under various adverse conditions across ACDC and DELIVER datasets, showing that Vireo consistently outperforms existing methods. (II) Qualitative visualization of segmentation predictions (left) and attention maps (right) under extreme night scenes, illustrating Vireo’s robustness and precise alignment with semantic cues. (III) Architectural comparison: (a) Traditional OVSS and (b) DGSS pipelines freeze or fine-tune VFM separately without cross-modal integration; (c) Our proposed Vireo introduces GeoText Query and Depth-VFM integration to enhance both semantic alignment and domain robustness.
@@
Moreover, in cross-domain scenarios, depth and geometric cues are largely insensitive to variations in illumination and texture[19, 20]. They supply reliable spatial constraints, easing the distribution shift of RGB features and sharpening boundary localization. Recent studies such as DepthForge [21] have shown that injecting depth query into a frozen VFM boosts domain generalization. Motivated by these findings, we adopt DepthAnything V2 [22] as our Depth VFM: its diverse pre-training delivers consistent depth estimation across domains, and keeping it fully frozen incurs minimal training cost while preserving real-time inference speed.
In this paper, we propose a VFM-based single-stage pipeline for OV-DGSS, termed Vireo. Specifically, in the encoder, both the VFM and DepthAnything modules are kept frozen. The VFM is leveraged to robustly capture cross-domain visual features, while DepthAnything extracts the scene’s intrinsic geometric structure. On this basis, we introduce Geometric-Text Query (GeoText Query) to fuse the extracted structural features with manually provided textual cues, progressively refining the feature maps between the frozen VFM layers. To mitigate the slow convergence caused by sparse gradients propagating through the frozen encoder, we introduce a Coarse Mask Prior Embedding (CMPE) at the beginning of the decoder to inject denser gradient signals. This design accelerates the convergence of mask supervision and further strengthens the influence of textual priors. Subsequently, we design the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH) to strengthen the synergistic integration of structural and textual modalities, ensuring that both cross-domain structural features and open-vocabulary cues learned by the GeoText Query are fully utilized in the final prediction.
For challenge 1: We find that GeoText Query not only capture structural and semantic cues within the frozen VFM encoder but also guide the progressive refinement of its feature representations. For challenge 2: The CMPE enhances gradient back-propagation into the encoder, while the redesigned DOV-VEH deepens the fusion of visual and textual priors. These components form a unified framework that simultaneously achieves domain robustness and strong open-vocabulary generalization. Our main contributions are summarized as follows:
@@
(2) We introduce GeoText Query to progressively refine frozen VFM features by injecting geometric cues from DepthAnything and aligning them with textual semantics, enabling structural-semantic fusion across encoder layers.
@@
Figure 2: Overview of the proposed Vireo framework for OV-DGSS. (a) Framework Overview. Geometric-Text Query permeate our model: they’re injected in Tunable Vireo to align domain priors, reused in CMPE to guide multi-scale feature fusion, and employed in DOV-VEH as queries for open-vocabulary segmentation, forming a unified end-to-end loop. (b) Tunable Vireo. GeoText Query are used to inject structural-textual priors across multiple layers. Depth-aware fusion and attention optimization are applied between intermediate frozen layers to progressively refine representations. (c) Coarse Mask Prior Embedding (CMPE). Multi-scale features from VFM are combined with GeoText Query to provide dense supervision and gradient signals for downstream modules. (d) Domain-Open-Vocabulary Head (DOV-VEH). Multi-level features are processed by a Pixel Decoder and Transformer Decoder to produce final predictions guided by open-vocabulary text queries.
@@ 3 Methodology - Vireo
We formulate the open-vocabulary DGSS problem as predicting a pixel-wise segmentation mask from an input image and a set of text labels. Given an image $I$ and a set of open-vocabulary class names $C$, our goal is to produce a mask $M$ where each pixel is classified according to one of the classes in $C$. As illustrated in Figure 2, our framework Vireo tackles this by leveraging frozen vision and text encoders with additional trainable prompt and depth modules for improved generalization, it consists of three key modules tailored for the OV-DGSS: (1) Tunable Vireo with GeoText Query: Introduces GeoText Query to inject and refine both geometric and textual information across layers of a frozen VFM. (2) Coarse Mask Prior Embedding (CMPE): Generates coarse prior masks to guide segmentation and reinforce gradient flow from the decoder back to the frozen encoder. (3) Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH): Integrates visual, geometric, and semantic features to produce final OV-DGSS predictions.
@@ 3.1 Overview
The input RGB image $I\in\mathbb{R}^{H\times W\times 3}$ is duplicated and independently fed into two frozen encoders: a VFM encoder $\mathcal{F}^{V}$ and a DepthAnything encoder $\mathcal{F}^{D}$. These backbones extract a series of multi-scale features $f^{V}_{l}$ and $f^{D}_{l}$ respectively, where $l$ denotes the encoder layer index in $[1,L]$. In parallel, a set of class labels $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ is transformed into natural language prompts $T=\{\texttt{prompt}_{1},\dots,\texttt{prompt}_{K}\}$, which are encoded using a frozen CLIP text encoder $\mathcal{F}^{T}$ to obtain textual embeddings $t_{k}\in\mathbb{R}^{d}$. These features serve as shared semantic priors across the entire framework and are precomputed once at initialization. In the Tunable Vireo module, each layer receives a tuple $(f^{V}_{l},f^{D}_{l},t_{k})$, and applies the GeoText Query $P_{l}$ to align and refine the visual features. This is achieved by computing cross-modal attention maps $\mathcal{A}_{l}=\text{Attn}(P_{l},f^{V}_{l},f^{D}_{l},t_{k})$ followed by feature fusion and projection. The refined output $\hat{f^{V}}_{l}$ is then forwarded back to update the VFM’s layer-wise activations.
@@
We select the refined visual features $\{{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder at layers $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$ for downstream decoding. In the Coarse Mask Prior Embedding (CMPE) module, these features are first upsampled to a unified spatial resolution and passed through a channel-spatial attention gating function $\mathcal{G}(\cdot)$. The gated outputs are then fused to form a global coarse feature $f^{M}\in\mathbb{R}^{d}$. The $f^{M}$ is projected and matched with the text embeddings $t_{k}\in\mathbb{R}^{d}$ to produce a coarse class probability map $\mathcal{M}\in\mathbb{R}^{H\times W\times K}$, where $\mathcal{M}$ serves both as a weak supervision signal and as a prior to construct the query embeddings $\mathcal{Q}$. The query prior is added to the GeoText Query and forwarded to the final segmentation head. In the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH), the multi-scale features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ are passed through a pixel decoder $\mathcal{D}_{p}(\cdot)$ to enhance spatial representation, followed by a Transformer decoder $\mathcal{D}_{T}(\cdot)$ that leverages positional embedding. The GeoText Query serve as learnable queries and interact with both the decoded features and the text embeddings, producing pixel-level mask embeddings $\mathcal{E}_{\text{mask}}(x,y)\in\mathbb{R}^{d}$ and classification embeddings $\mathcal{E}_{\text{cls}}(k)\in\mathbb{R}^{d}$. The final prediction is $\hat{\mathcal{M}}(x,y,k)\in\mathbb{R}^{H\times W\times K}$ provides the pixel-wise semantic prediction with both fine-grained detail and open-vocabulary generalization capability.
3.2 Tunable Vireo with GeoText Query
@@
During inference, the input image $I\in\mathbb{R}^{H\times W\times 3}$ is simultaneously processed by a frozen visual encoder $\mathcal{F}^{V}$ and a frozen depth encoder $\mathcal{F}^{D}$ (e.g., DepthAnything). For each selected layer $l\in\{1,\dots,L\}$, we obtain the visual feature map $f^{V}_{l}=\mathcal{F}^{V}_{l}(I)$ and the depth feature map $f^{D}_{l}=\mathcal{F}^{D}_{l}(I)$. Each layer of the Tunable Vireo module receives the tuple $(f^{V}_{l},f^{D}_{l},\{t_{k}\})$ along with a layer-specific GeoText Prompt $P_{l}\in\mathbb{R}^{N\times d}$, where $N$ is the number of learnable queries.
@@ 3.3 Coarse Mask Prior Embedding (CMPE)
Each $f^{\text{class}}_{k}$ is projected into the embedding space as $e^{\text{class}}_{k}\in\mathbb{R}^{d}$, and combined with a set of learnable queries vectors $\{q_{j}\}_{j=1}^{N_{q}}$ to produce the query priors:
@@
The final query priors $\{q^{\text{prior}}_{j}\}_{j=1}^{N_{q}}$ are added to the corresponding GeoText Query and forwarded into the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH).
@@ 3.4 Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH)
The DOV-VEH module receives the refined multi-scale features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder, where $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$, along with the updated GeoText Query $\{P_{l}\}$. These features are first processed by a pixel decoder $\mathcal{D}_{p}(\cdot)$, which leverages multi-scale cross-attention to extract rich spatial context: $f^{\text{pix}}=\mathcal{D}_{p}\left(\hat{f}^{V}_{l_{i}}\right)$. The fused feature $f^{\text{pix}}\in\mathbb{R}^{H\times W\times d}$ is then compressed via a $1\times 1$ convolution and enriched with sinusoidal positional encoding to preserve spatial structure.
The enhanced features are fed into a Transformer Decoder $\mathcal{D}_{T}(\cdot)$, where the GeoText Query act as learnable queries. Through stacked layers of self-attention and cross-attention with $f^{\text{pix}}$, the model captures fine-grained visual-semantic alignment, yielding a set of high-resolution mask features $\mathcal{E}_{\text{mask}}(x,y)\in\mathbb{R}^{d}$ at each spatial position.
Simultaneously, the GeoText Query $\{P\}$ are passed through a two-layer MLP and then interact with the text embeddings $\{t_{k}\}$ (precomputed from the CLIP text encoder) to produce classification-level representations $\mathcal{E}_{\text{cls}}(k)\in\mathbb{R}^{d}$. The final segmentation prediction $\hat{\mathcal{M}}\in\mathbb{R}^{H\times W\times K}$ is generated via an Einstein summation over the two embeddings:
$\hat{\mathcal{M}}(x,y,k)=\sum_{d=1}^{D}\mathcal{E}_{\text{mask}}(x,y,d)\cdot\mathcal{E}_{\text{cls}}(k,d),$ (5)
@@ 4.4 Ablation Study
Figure 5: Visualizations of attention maps and affinity maps under differnet scenes, where CMPE denotes the Coarse Mask Prior Embedding, GTQ represents the GeoText Query, respectively.
@@
++ concat $f^{d}_{i}$ 70.8 56.1 79.4 72.8 63.6 74.2
++ Prompt Depth Anything 70.4 55.5 79.7 72.0 63.8 73.9
++ Depth Anything V2 71.5 56.7 80.5 73.3 64.4 74.5
++ DA + AO 72.2 57.4 80.9 74.2 65.1 75.0
++ DOV-VEH 70.9 56.2 79.8 72.8 63.7 74.2
++ CMPE 71.6 56.9 80.2 73.4 64.1 74.6
++ GeoText Query 74.0 58.4 81.1 74.8 65.3 75.3
@@
Fine-tune
Trainable
Params
mIoU Avg.
@@
Fine-tune
Trainable
Params
mIoU Avg.
@@
Fine-tune
Trainable
Params
mIoU Avg.
@@
Fine-tune
Trainable
Params
mIoU Avg.
@@
In Table 6, using Depth Anything V2 alone yields an approximate $0.6\%$ mIoU gain across all six scenarios, and adding depth augmentation with attention optimization (DA + AO) delivers a further $\approx 1\%$ improvement. The DOV-VEH and CMPE modules each add $0.5\%$$0.8\%$, validating mask vectors and dense gradient embedding. GeoText Query on its own provides a substantial $\approx 4.4\%$ boost, underscoring the complementary benefits of fusing semantic and geometric cues.
Pronounced Attention Focus. Figure 5 further confirm that GeoText Query steer the model toward geometry-sensitive regions, while CMPE strengthens gradients on foreground masks; compared with baselines, Vireo’s focus is tighter on scene structure and semantic boundaries, explaining its consistent advantage in cross-domain and open-vocabulary settings.